OpenStudy (kainui):

What is the coefficient on x^(n-1) on the expansion of (x+1)(x+2)(x+3)...(x+n) ?

OpenStudy (anonymous):

you want a guess?

OpenStudy (anonymous):

my guess is \(n!\) but i could easily be wrong experiment and see what it looks like for \(n=3,4,5\)

OpenStudy (kainui):

Actually n! is the coefficient on the x^0 term =)

OpenStudy (anonymous):

yeah my answer was way off

OpenStudy (loser66):

the sequence is 1,3, 6, 10,.....

OpenStudy (anonymous):

is the next one 15?

OpenStudy (loser66):

I mean: (x+1)(x+2) , the coefficient of x^1 = 1 (x+1)(x+2)(x+3) , the coefficient of x^2 is 3

OpenStudy (loser66):

and so on...

OpenStudy (anonymous):

yes it is , and then 21

OpenStudy (anonymous):

sequence 1,3,6,10,15, 21 would mean you have a quadratic

OpenStudy (anonymous):

you can tell it is quadratic because the differences go up by one each time

OpenStudy (anonymous):

i could be wrong, but it might be \(\frac{1}{2}(n^2+n)\) try that

OpenStudy (anonymous):

better guess than \(n!\) for sure!

OpenStudy (kainui):

Yeah that's correct =) Now I dare someone to try to find the x^(n-2) or x^(1) coefficients, or the general formula for any coefficient. =)

OpenStudy (anonymous):

oh, i didn't realize this was a test

OpenStudy (kainui):

Oh yeah just a fun game I'm playing haha

OpenStudy (anonymous):

do you know the answer?

OpenStudy (anonymous):

and it is it a quadratic as well?

OpenStudy (anonymous):

i meant "polynomial" not quadratic

OpenStudy (kainui):

Nope, it's not quadratic as far as I know. For instance, the x^1 term's coefficient is \[\Large n \sum_{k=1}^n \frac{1}{k}\] But I don't know the closed form of this, if there even is one. The weird part is I know the form of all the higher terms but I have no idea how to even express them with summation signs.

OpenStudy (anonymous):

What do you know about Stirling numbers?

OpenStudy (kainui):

Huh, no I have never heard of these before, but this is quite interesting.

OpenStudy (kainui):

Here is my sort of insight that I don't know how to write: \[\Large f(x) = (x+a)(x+b)(x+c) \\ f(x) = abc[ (\frac{1}{abc})x^3 + (\frac{1}{ab}+ \frac{1}{ac}+\frac{1}{bc})x^2+(\frac{1}{a}+\frac{1}{b}+\frac{1}{c})x + 1]\] So each coefficient on the x^k term is really just a sum of fractions of all the possible ways to pick k of the roots and multiply them together.